Project

General

Profile

NDOS Commissioning Meeting 2011/10/06

Attendance: Jon, Mark Mathis, Luke, Mat, Leon, Dave Demuth, Susan, Nick, Zeb, Rafael, Jan, Enrique, Martin, Andre, Tyler, Mike, Kurt, Andrew, Peter

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & Detector Update

  • Beam:
  • Detector:
    • APDs moved last week:
      • shuffled 15 APDs from block 0 to blocks 1 and 2
      • 13/15 are operating well, basically all pixels are reading out that were before (~4 APDs have two or fewer pixels masked off)
      • 1 APD was found to have trim pot pulled off; attempted repair shows HV looks right, but APD is noisy.
      • 1 APD is exhibiting the usual surface contamination problem. Leon will take this back to CalTech
    • APD work this week:
      • Leon and Mat installed a silicone-coated APD with intentional leak (o-ring was left out). That APD was noisy last week, but after cable-ectomy it is behaving well (for the last few days)
      • installed a parylene-coated APD with intentional leak (o-ring was left out). That is operating normally.
      • Leon brought out vacuum-testing fixtures; we also received a shipment of sealed heatsinks, so these are being vacuum tested and so far 9 have been installed, 3 of which have parylene-coated APDs. Of these 9, one is showing noisy behavior.
      • We will enable cooling on 7 of these APDs after this meeting, which will include the two leaky APDs, 3 units with new heatsinks but no coating, and 2 parylene-coated APDs with new heatsinks.
      • if all goes well tonight, then we'll enable cooling on the remaining three APDs on the detector.
      • DSO scans will be done once per shift through this weekend.
    • Mat: will also be working with David Demuth tomorrow in order to test assembly tracking.

DCS Update

  • Martin:
    • Zukai's "event display" overview is now in use
    • GUI distinguishes between good warm and good cold APDs (dark vs. light green)
    • on-call experts (24/7) are now listed on the white-board as well as the DCS wiki.

DAQ Update

  • Peter: working on updating (cleaning up) the base release of the DAQ software
  • Jon: RC has been updated, and instructions for shifters has changed.
  • Andrew: new production TDU out at MINOS building, and a new antennae has been installed that should improve stability, but it's not in use yet.

AOB